# Geometry-Informed Neural Networks

### v3

@@ D.1 Loss plots
In Figures 9 and 10 we show the loss plots for training single and multiple shapes respectively. As expected the unweighted losses (middle rows in the Figures) decrease, while the Lagrange terms (bottom rows) increase over training.
Figure 9: Loss plots for training a single shape. The lines with higher alpha are exponential-moved averages of the lower-alpha values by the factor 0.99. (a) The top plot shows the losses as used for backpropagation. (b) The middle plot shows the unweighted losses for individual constraints. (c) The bottom plots show the $\lambda$ values for the individual constraints.
Figure 10: Loss plots for training multiple shapes. The lines with lower transparency are exponential-moved averages with factor 0.99 of the higher-transparency values. (a) The top plot shows the losses as used for backpropagation. (b) The middle plot shows the unweighted losses for individual constraints. (c) The bottom plots show the $\lambda$ values for the individual constraints.
Appendix E Connectedness
We provide additional details on our approach to the connectedness loss. We break this down in three parts: First, we define the signed distance function of a shape $\Omega$ which the neural field we train approximates. Then, we give a short rundown on computing the persistent homology (PH), in particular the PH of a neural field in a not rectangular region. Lastly, we explain how to obtain a differentiable loss on the field from the outputs of the, in general non-differentiable, PH computation.
@@ Appendix E Connectedness
A point $x\in X$ belongs to the medial axis if its closest boundary point is not unique. The gradient of an SDF obeys the eikonal equation $\|\nabla f(x)\|=1$ everywhere except on the medial axis where the gradient is not defined. In INS, the SDF is approximated by a NN with parameters $\theta$: $f_{\theta}\approx f$.
Connectedness refers to an object $\Omega$ consisting of a single connected component. It is a ubiquitous feature enabling the propagation of mechanical forces, signals, energy, and other resources. Consequentially, connectedness is an important constraint for enabling GINNs. In the context of machine learning, connectedness constraints have been multiply applied in segmentation [78, 79, 80], surface reconstruction [81], and 3D shape generation with voxels [82], point-clouds [83] and INSs [84].
@@ E.1 Persistent Homology
Persistent Homology is one of the primary tools which has emerged from topological data analysis (TDA) to extract topological features from data. Data modalities such as point clouds, time series, graphs and $n$-dimensional images can all be transformed into weighted cell complexes from which the homology can be computed. The homology provides global information about the underlying data and is generally robust.
@@
$\varnothing\subseteq X_{1}\subseteq\dots\subseteq X_{N}=X$ (23)
@@
Persistent Homology encodes how the homology of an increasing sequence of complexes changes under a given filtration. Topological features appear and and vanish as the filter function sweeps over $X$. The birth time $b$ of a feature defined as the value $\alpha_{n}$ at which the homology of $C_{n}(X_{n})$ changes to include this feature. The death time $d$ of a feature is analogously defined as the value $\alpha_{n}$ at which it is removed from $C_{n}(X_{n})$. The persistence of a feature is defined as the length of its lifetime $l=d-b$.
For each Betti number $b_{n}$ (for each homology class $H_{n}$) the information about the persistent homology of a given filtration is encoded in a persistence diagram containing the points $(b,d)$ of the birth and death pairs of all $n$-dimensional topological features (changes in the dimension of $H_{n}$). The persistence diagrams contain the entire topological information about underlying the space or shape for a sufficiently fine filtration.
To compute the persistent homology of a neural field, we evaluate the network on a cubical complex on the domain of the field, i.e. a grid in $\mathbb{R}^{N}$. The output is simply a gray scale image (we are only dealing with scalar fields in this work) and the PH can computed with existing algorithms. The current SotA algorithm for PH computation on cubical complexes is CRipser [85].
@@
In general, we are interested in computing the PH within a given envelope, which is not necessarily a rectangular region. We achieve this by sampling the field in a rectangular domain containing the envelope and setting the value of points not in the envelope to $\infty$. Applying the PH computation to this altered image then correctly returns the evolution of persistence features within the envelope. The only drawback of this method is the additional computational cost of having to include the grid points outside the envelope in the PH computation, which is why the bounding domain should be chosen tightly around the envelope.
The PH computation itself does not have to be differentiable (and the CRipser library we use is not) because the cells, i.e. the grid point of the image, at which a given persistence feature is born or killed are stored. Hence we can simply use the network output at this grid coordinate to compute the loss and there are no issues concerning differentiability or having to re-implement the PH computation into PyTorch.
@@ E.2 Differentiable topology loss
Remark: The representatives of homology class are not uniquely determined. The CRipser library internally chooses a representative and then outputs its coordinates. In practice this caused no issues.
@@
More generally, to obtain a shape with a Betti number $b_{n}=m$ at iso-level $a_{n}$, the summation above runs from $i=m+1\dots M$. The full topology loss for a $N$-dimensional shape is then given as
@@ Appendix F Diversity
Figure 11: A visual comparison of different diversity losses in a simple 2D example ($\mathcal{F}=\mathbb{R}^{2}$ and the feasible set $\mathcal{K}$ is the partial annulus). Each point $f\in\mathcal{F}$ represents a candidate solution. The points are optimized to maximize the diversity within the feasible set. The top row shows the minimal aggregation $\delta_{\text{min}}$ as defined in Equation 26. The bottom row shows the total aggregation $\delta_{\text{sum}}$ as defined in Equation 27. Each column uses a different exponent $p\in\{0.5,1,2\}$. For $0\leqslant p\leqslant 1$ the minimal aggregation diversity $\delta_{\text{min}}$ is concave meaning it favors increasing smaller distances over larger distances. This leads to a uniform coverage of the feasible set. In contrast, the $\delta_{\text{min}}$ is convex for $p\geq 1$ as indicated by the formed clusters for $p=2$. Meanwhile, $\delta_{\text{sum}}$ pushes the points to the boundary of the feasible set for all $p$.
@@
Using different exponents $p\in\{1/2,1,2\}$ illustrates how $\delta_{\text{min}}$ covers the domain uniformly for $0\leqslant p\leqslant 1$, while clusters form for $p>1$. The total aggregation measure always pushes the samples to the extremes of the domain.

### v4

@@ D.1 Loss plots
In Figures 11 and 12, we show the loss plots for training single and multiple shapes, respectively. As expected, the unweighted losses (middle rows in the Figures) decrease, while the Lagrange terms (bottom rows) increase over training.
Figure 11: Loss plots for training a single shape of the bracket problem. The solid lines are exponential-moving averages (factor 0.99) of the noisy values in lighter colors. (a) The losses used for backpropagation. (b) The unweighted losses of each constraint. (c) The $\lambda$ values of each constraint.
Figure 12: Loss plots for training multiple shapes of the bracket problem. The solid lines are exponential-moving averages (factor 0.99) of the noisy values in lighter colors. (a) The losses used for backpropagation. (b) The unweighted losses of each constraint. (c) The $\lambda$ values of each constraint.
Appendix E Topology loss
We provide additional details on our approach to the connectedness loss. We break this down in three parts: First, we define the signed distance function of a shape $\Omega$, which the neural field we train approximates. Then, we give a short rundown on computing the persistent homology (PH), in particular the PH of a neural field in a non-rectangular region. Lastly, we explain how to obtain a differentiable loss on the field from the outputs of the non-differentiable PH computation.
@@ Appendix E Topology loss
A point $x\in X$ belongs to the medial axis if its closest boundary point is not unique. The gradient of an SDF obeys the eikonal equation $\|\nabla f(x)\|=1$ everywhere except on the medial axis, where the gradient is not defined. In INS, the SDF is approximated by a NN with parameters $\theta$: $f_{\theta}\approx f$.
Connectedness refers to an object $\Omega$ consisting of a single connected component. It is a ubiquitous feature enabling the propagation of mechanical forces, signals, energy, and other resources. Consequently, connectedness is an important constraint for enabling GINNs. In the context of machine learning, connectedness constraints have been multiply applied in segmentation (Wang et al., 2020; Clough et al., 2022; Hu et al., 2019), surface reconstruction (Brüel-Gabrielsson et al., 2020), and 3D shape generation with voxels (Nadimpalli et al., 2023), point clouds (Gabrielsson et al., 2020), and INSs (Mezghanni et al., 2021).
@@ E.1 Persistent Homology
Persistent homology (PH) is one of the primary tools that has emerged from topological data analysis to extract topological features from data. Data modalities such as point clouds, time series, graphs, and $n$-dimensional images can all be transformed into weighted cell complexes from which the homology can be computed. The homology provides global information about the underlying data and is generally robust.
@@
$\varnothing\subseteq X_{1}\subseteq\dots\subseteq X_{N}=X\ .$ (23)
@@
Persistent Homology encodes how the homology of an increasing sequence of complexes changes under a given filtration. Topological features appear and vanish as the filter function sweeps over $X$. The birth time $b$ of a feature is defined as the value $\alpha_{n}$ at which the homology of $C_{n}(X_{n})$ changes to include this feature. The death time $d$ of a feature is analogously defined as the value $\alpha_{n}$ at which it is removed from $C_{n}(X_{n})$. The persistence of a feature is defined as the length of its lifetime $l=d-b$.
For each Betti number $b_{n}$ (for each homology class $H_{n}$) the information about the persistent homology of a given filtration is encoded in a persistence diagram containing the points $(b,d)$ of the birth and death pairs of all $n$-dimensional topological features (changes in the dimension of $H_{n}$). For a sufficiently fine filtration, the persistence diagrams contain the entire topological information about the underlying space or shape.
To compute the persistent homology of a neural field, we evaluate the network on a cubical complex on the domain of the field, i.e., a grid in $\mathbb{R}^{N}$. The output is simply a gray-scale image (since we are only dealing with scalar fields in this work), and the PH can be computed with existing algorithms. The current state-of-the-art algorithm for PH computation on cubical complexes is CRipser (Kaji et al., 2020).
@@
In general, we are interested in computing the PH within a given design region or envelope, which is not necessarily a rectangular region. We achieve this by sampling the field in a rectangular domain containing the envelope and setting the value of points not in the envelope to $\infty$. Applying the PH computation to this altered image then correctly returns the evolution of persistence features within the envelope. The only drawback of this method is the additional computational cost of having to include the grid points outside the envelope in the PH computation, which is why the bounding domain should be chosen tightly around the envelope.
The PH computation itself does not have to be differentiable (and the CRipser library we use is not) because the cells, i.e., the grid points of the image, at which a given persistence feature is born or killed, are stored. Hence, we can simply use the network output at this grid coordinate to compute the loss and there are no issues concerning differentiability or having to re-implement the PH computation into PyTorch.
@@ E.2 Differentiable topology loss
Remark: The representatives of a homology class are not uniquely determined. The CRipser library internally chooses a representative and then outputs its coordinates. In practice, this caused no issues.
@@
More generally, to obtain a shape with a Betti number $b_{n}=m$ at iso-level $a_{n}$, the summation above runs from $i=m+1\dots M$. The full topology loss for an $N$-dimensional shape is then given as
@@ Appendix F Diversity
Figure 13: A visual comparison of different diversity losses in a simple 2D example ($\mathcal{F}=\mathbb{R}^{2}$ and the feasible set $\mathcal{K}$ is the partial annulus). Each point $f\in\mathcal{F}$ represents a candidate solution. The points are optimized to maximize the diversity within the feasible set. The top row shows the minimal aggregation $\delta_{\text{min}}$ as defined in Equation 26. The bottom row shows the total aggregation $\delta_{\text{sum}}$ as defined in Equation 27. Each column uses a different exponent $p\in\{0.5,1,2\}$. For $0\leq p\leq 1$ the minimal aggregation diversity $\delta_{\text{min}}$ is concave meaning it favors increasing smaller distances over larger distances. This leads to a uniform coverage of the feasible set. In contrast, the $\delta_{\text{min}}$ is convex for $p\geq 1$ as indicated by the formed clusters for $p=2$. Meanwhile, $\delta_{\text{sum}}$ pushes the points to the boundary of the feasible set for all $p$.
@@
Using different exponents $p\in\{1/2,1,2\}$ illustrates how $\delta_{\text{min}}$ covers the domain uniformly for $0\leq p\leq 1$, while clusters form for $p>1$. The total aggregation measure always pushes the samples to the extremes of the domain.